OpenStudy (ksaimouli):

True or false r(x)=2x-1/(x=2)(x-1) changes sign at x=-2

OpenStudy (ksaimouli):

help me bro

OpenStudy (anonymous):

r(x)=2x-1/(x=2)(x-1) = what is this

OpenStudy (ksaimouli):